identify the conic section if it is a parabola give the vertex if it is a circle give the center and radius if it is a ellipse or a hyperbola give the center and foci
x^2+y^2-4x+6y=0

Answer by Boreal(15235)

y^2+6y+x^2-4x=0
y^2+6x+9+x^2-4x+4=13, completing the square.
(y+3)^2 +(x-2)^2=13
Circle with center at (2,-3) and radius square root (13)